OpenStudy (anonymous):

Which expression is equivalent to 46 ⋅ 4−8?

OpenStudy (acxbox22):

using PEMDAS multiply 46 by 4 first then subtract 8 tell me what you get?

OpenStudy (anonymous):

176 @acxbox22

OpenStudy (acxbox22):

yes, good job

OpenStudy (anonymous):

But that is none of the answers in my answer choices @acxbox22

Directrix (directrix):

Could this be: 46 ⋅ 4 − 8 = 46 * 4 and then subtract 8 46 ⋅ 4 − 8 = 46 * 4 - 8 46 ⋅ 4 − 8 = 184 - 8 = ? @timeflybyegirl